OpenStudy (anonymous):

Given a central angle of 75 degrees, find the length of its intercepted arc in a circle of radius 15 centimeters.

OpenStudy (anonymous):

\[25\pi/4\]the whole chircle would be \[2*\pi*15\] and would have a central angle of 360 degrees A central angle of 75 degrees in teh same circle would intersect an arc lenght of \[(75/360)*2*\pi*15 = 25\pi/4\]